The Y arm cavity length was locked to the PSL frequency with ~26kHz UGF, and 25 degrees phase margin. Slow actuation was done on ETMY using CM_Slow as an error signal, while fast actuation was done on the IMC error point via the IN2 input of the IMC servo board. Attachment #1 shows the comparison of the in-loop error signal spectra with only slow actuation and with the full CM loop engaged.
I hypothesize that the high-frequency noise (>100 Hz) is higher for POY than POX in Attachment #1 because I am using the "MON" port of the demod board - this has a gain of 2, and there could also be some flaky components in this path, hence the high frequency noise is a factor of a few greater in the POY spectrum relative to the POX spectrum (which is using the main demodulated output). For REFL11, we have a low noise preamp generating the input signal so I don't think we need to worry about this too much.
The PC Drive RMS didn't look any stranger than it usually does for the duration of the lock.
Attachment #2 shows the OLTF of the locking servo with the final gains / settings, which are in bold. The loop is maybe a bit marginal, could possibly benefit from backing off one of the super boosts. But the arm has stayed locked for >1 hour.
The purpose of this test was to verify the functionality of the CM board and also the IN2 of the IMC servo board in a low-pressure environment. Once I confirm that the modelled OLTF lines up with the measured, I will call this test a success, and move on to looking at REFL11 in the arms on ALS, PRMI on 3f config. I am returning the REFL11 signal to the input of the CM board, but the SR785 remains hooked up.
Unrelated to this work - PMC alignment was tweaked to improve input power to IMC by ~5%.
One of the differences between the direct POY and the CM_SLOW POY is the presence of the CM Servo gain stages. So this might mean that you need to move some of the whitening gain to the CM IN1 gain.
In preparation for trying out some high-bandwidth Y arm cavity locking using the CM board, I hooked up the POY11_Q_Mon channel of the POY11 demod board to the IN1 of the CM board (and disconnected the usual REFL11 cable that goes to IN1). The digital phase rotation for usual POY Yarm locking is 106 degrees, so the analog POY11_Q channel contains most of the signal. I then set the IN1 gain of the servo to 0dB, and looked at the CM_Slow signal - I changed the whitening gain of this channel to +18dB (to match that used for POY11_I and POY11_Q), and found that I had to apply a digital gain of 0.5 to get the PDH horns in the usual POY11_I signal and the CM_Slow signal to line up. There was also a sign inversion. Then I was able to use the digital LSC system and lock the Y arm cavity length to the PSL frequency by actuating on ETMY, using CM_Slow as an error signal. A comparison of the in-loop POY11_I ASD when the arm is locked is shown in Attachment #1 - CMslow seems to be dominated by some kind of electroncis noise above ~100 Hz, so possibly needs more whitening (even though the nominal whitening filter is engaged)?
Anyway, now that I have this part of the servo working, the next step is to try and engage the AO path and achieve a higher BW lock of the Y arm cavity to the PSL frequency (= IMC length). Maybe it makes more sense to actuate on MC2 for the slow path.
After the QPD fix, both arms report consistent buildup - see Attachment #1. The peak values touch ~250, corresponding to a PRG of ~13. The IFO becomes critically coupled at PRG=15. I am finding that the 3f signal offsets are changing as a function of the CARM offset, and this could be responsible for the lock breaking as I approach 0 CARM offset. I found that I could maintain a more stable and deterministic transition to zero CARM offset by dynamically adjusting the 3f PRCL error signal offset to keep the REFL11 signal approximately at 0. Some shaking seems to have commenced so I am breaking for now.
Note that I find scattered throughout the elog references to a similar problem of the PRMI losing lock as the CARM offset is reduced, e.g. here. But haven't stumbled across what the resolution was, the PRFPMI could be locked pretty easily in 2015 I remember.
Koji and I had noticed that there was some discrepancy between the switchable gain stages of the EX and EY QPDs. Sadly, there was no indication that these switches even exist on the QPD MEDM screen. Yehonathan and I rectified this today. Both EX and EY Transmon QPDs now have some extra info (see Attachment #1). We physically verified the indicated quadrant mapping for the EX QPD (see previous elog in this thread for the details), and I edited the screen accordingly. EY QPD still has to be checked. Note also that there is an ND=0.4 + ND=0.2 filter and some kind of 1064nm light filter installed in series on the EX QPD. The ND filters have a net transmission T~25%.
After making the EX and EY QPDs have the same switchable gain settings (I also reset the trans normalization gains), the angular motion witnessed is much more consistent between the two now - see Attachment #2. The high-frequency noise of the sum channel is somewhat higher for EX than EY - maybe the ND filters are different on the two ends?
Note that there was an extra factor of 40 gain on the EX QPD relative to EY during the lock yesterday. So really, the signals were probably just getting saturated. Now that the gains are consistent between the ends, it'll be interesting to see how balanced the buildup of the two arms is. There still remains the problem that the MICH loop was too unstable, which probably led to to excess arm power fluctuations.
There is a mark on the QPD surface that is probably dirt (since we never have such high power transmitted through the ETM to damage the QPD). I'll try cleaning it up at the next opportunity.
In search of the source of discrepancy between the QPD readings in the X and Y arms, we look into the schematics of the QPD amplifier - DCC #D990272.
We find that there are 4 gain switches with the following gain characteristics (The 40m QPD whitening board has an additional gain of 4.5):
Switch 4 bypasses the amps controlled by switch 2 and 3 when it is set to 1 so they don't matter in this state.
Note that according to elog-13965 the switches are controlled through the QPD whitening board by a XT1111a Acromag whose normal state is 1.
Also, according to the QPD amplifier schematics, the resistor on the transimpedance, controlled by switch 1, is 25kOhm. However, according to the EPICS it is actually 5kOhm. We verify this by shining the QPD with uniform light from a flashlight and switching switch1 on and off while measuring the voltages of the different segments. The schematics should be updated on the DCC.
Surprisingly, QPDX switches where 0,0,0,0 while QPDY switches where 1,0,0,1. This explains the difference in their responses.
We check by shining a laser pointer with a known power on the different segments of QPDX that we get the expected number of counts on the ADC and that the response of the different segments is equal.
I took a look at the TRX/TRY RIN reported in the single arm POX/POY lock, and compared the performance of the two available PDs at each of the two ends. Attachment #1 shows the results. Some remarks:
The shaking started earlier today than yesterday, at ~9pm local time.
While the IFO is shaking, I thought (as Jan Harms suggested) I'd take a look at the cross-spectra between our seismometer channels at the dominant excitation frequency, which is ~1.135 Hz. Attachment #1 shows the phase of the cross spectrum taken for 10 averages (with 30mHz resolution) during the time period when the shaking was strong yesterday (~1500 seconds with 50% overlap). The logic is that we can use the relative phasing between the seismometer channels to estimate the direction of arrival and hence, the source location. However, I already see some inconsistencies - for example, the relative phase between BS_Z and EX_Z suggests that the signal arrives at the EX seismometer first. But the phasing between EX_Y and BS_Y suggests the opposite. So maybe my thinking about the problem as 3 co-located sensors measuring plane-wave disturbances originating from the same place is too simplistic? Moreover, Koji points out that for two sensors separated by ~40m, for a ground wave velocity of 1.5 km/s, the maximum phase delay we should see between sensors is 30 msec, which corresponds to ~10 degrees of phase. I guess we have to undo the effects of the phasing in the electronics chain.
Does anyone have some code that's already attempted something similar that I can put the data through? I'd like to not get sucked into writing fresh code.
🤞 this means that the shaking is over for today and I get a few hours of locking time later today evening.
Another observarion is that even after the main 1.14 Hz peak dies out, there is elevated seismic acitivity reported by the 1-3 Hz BLRMS band. This unfortunately coincides with some stack resonance, and so the arm cavity transmission reports greater RIN even after the main peak dies out. Today, it seems that all the BLRMS return to their "nominal" nighttime levels ~10 mins after the main 1.14 Hz peak dies out.
Jon and I were surveying the CDS situation so that he can prepare a report for discussion with Rolf/Rich about our upcoming BHD upgrade. In our poking around, we must have bumped something somewhere because the c1ioo machine went offline, and consequently, took all the vertex models out. I rebooted everything with the reboot script, everything seems to have come back smoothly. I took this opportunity to install some saturation counters for the arm servos, as we have for the CARM/DARM loops, because I want to use these for a watch script that catches when the ALS loses lock and shuts stuff off before kicking optics around needlessly. See Attachment #1 for my changes.
Some ideas Koji suggested:
For the second idea, it is convenient to be able to control the arms in the XARM/YARM basis as opposed to the CARM/DARM basis as we usually do when going through the locking sequence. After some fiddling, I am able to reliably execute this transition, and achieve a state where the FP arm cavities are resonant for the IR with the ALS beat note frequency being the error signal being used. Some important differences:
the upgrade seems to have been successfully executed - the machine was restarted at ~430pm local time. Projector remains off and diagnostic striptools are on the samsung.
and so it begins...until this is finished I have turned off the projector and moved the striptools to the big TV (time to look for Black Friday deals to replace the projector with a 120 inch LED TV)
The nightly seismic activity enhancement continued during the weekend. It always shows up around 10pm local time, persists for ~1 hour, and then goes away. This isn't a show stopper as long as it stops at some point, but it is annoying that it is eating up >1 hour of possible locking time. I walked over to CES, no one there admitted to anything - there is an "Earth Surface Dynamics Laboratory" there that runs some heavy equipment right next to us, but they claim they aren't running anything post ~530pm. Rick (building manager ?) also doesn't know of anything that turns on with the periodicity we see. He suggested contacting Watson but I have no idea who to talk to there who has an overview of what goes on in the building. 😢
Attachment #1 is a spectrogram of the BS sesimometer signals for a ~24 hour period (from Wednesday night to Thursday night local time, zipped because its a large file). I've marked the nearly pure tones that show up for some time and then turn off. We need to get to the bottom of this and ideally stop it from happening at night because it is eating ~1 hour of lockable time.
We considered if we could look at the phasing between the vertex and end seismometers to localize the source of the disturbance.
The DC port of the Bias-Tee is routed to (a modified version of) the iLIGO whitening board. This has the well-known problem of the protection diodes of the LT1125 quad-op-amp lowering the (ideally infinite) input impedance of the first gain stage (+24 dB). To be sure as to how much signal we can put into this port (in anticipation of trying some variable finesse PRFPMI locking but also for general book-keeping), I tested the usable input range by driving a triangle wave at ~3 Hz and changing the amplitude of the signal until we observed saturation. We found that we could drive a 10 Vpp signal at which point there was evidence of some clipping (it was asymmetric, the top end of the signal was getting clipped at +14,000 cts while the bottom end still looked like a triangle wave at -16,000 counts). Anyway we probably don't want to exceed +/- 10,000 counts on this channel. This is consistent with Hartmut's statement of having +/- 4V of usable range (although the counts he mentions are twice what I saw yesterday).
Other discussion points between Rana, Koji and Gautam:
The clue was that the loop X arm POX loop looked to have low (<3dB)) gain margin around 600 Hz (and again at 700 Hz). Attachment #1 shows a comparison of the OLTF for this loop (measured using the IN1/IN2 method) before and after our change. We hypothesize that one of the violin filters that were turned off had non-unity DC gain, because I had to lower the loop gain by 20% after these turn-offs to have the same UGF. I updated the snap files called by the arm locking scripts.
I think I caught all the places where the FM settings are saved, but some locking scripts may still try and turn on some of these filters, so let's keep an eye on it.
We turned off many excessive violin mode bandstop filters in the LSC.
The large ground motion at 1 Hz started up again tonight at around 23:30. I walked around the lab and nearby buildings with a flashlight and couldn't find anything whumping. The noise is very sinusoidal and seems like it must be a 1 Hz motor rather than any natural disturbance or traffic, etc. Suspect that it is a pump in the nearby CES building which is waking up and running to fill up some liquid level. Will check out in the morning.
Estimate of displacement noise based on the observed MC_F channel showing a 25 MHz peak-peak excursion for the laser:
dL = 25e6 * (13 m / (c / lambda)
= 1 micron
So this is a lot. Probably our pendulum is amplifying the ground motion by 10x, so I suspect a ground noise of ~0.1 micron peak-peak.
(this is a native PDF export using qtgrace rather than XMgrace. uninstall xmgrace and symlink to qtgrace.)
Due to some feedforward work by Jenne or EQ some years ago, we have had ~10 violin notches on in the LSC between the output matrix and the outputs to the SUS.
They were eating phase, computation time, and making ~3 dB gain peaking in places where we can't afford it. I have turned them off and Gautam SDF safed it.
Offensive BS shown in brown and cooler BS shown in blue.
To rotate the DTT landscape plot to not be sideways, use this command (note that the string is 1east, not least):
pdftk in.pdf cat 1east output out.pdf
at 1 Hz' this effect is not large so that's real translation. at lower frequencies a ground tilt couples to the horizontal sensors at first order and so the apparent signal is amplified by the double integral. drawing a free body diagram u can c that
x_apparent = (g / s^2) * theta
but for vortical this not tru because it already measures the full free fall and the tilt only shows up at 2nd order
Here is some disturbance in the spacetime curvature, where the local gradient of the metric seems to have been modulated (in the "downward" as well as in the other two orthogonal Cartesian directions) at ~1 Hz - seems real as far as I can tell, all the suspensions were being shaken about and all the seismometers witnessed it, though the peak is pretty narrow. A broader, less prominent peak also shows up around 0.5 Hz. We couldn't identify any clear source (no LN2 fill-up / obvious CES activity). This event lasted for ~45 mins, and stopped around 2315 local time. Shortly (~5min) after the ~1 Hz peak died down, however, the 3-10 Hz BLRMS channel reports an increase by ~factor of 2.
Onto trying some locking now that the suspensions have settled down somewhat.
this is due to the Equivalence Principle: local accelerations are indistinguishable from spacetime curvature. On a spherical Earth, the local gradient of the metric points in the direction towards the center of the Earth, which is colloquially known as "down".
I don't understand why the z-axis motion reported by the T240 is ~10x lower at 10 mHz compared to the X and Y motions. Is this some electronics noise artefact?
Attachment #1 shows the spectra of our three available seismometers over a period of ~10ksec.
Attachment #2 shows the result of applying frequency domain Wiener filter subtraction to the POP QPD (target) with the vertex seismometer signals as witness channels.
Today I realized that pip and other python2,3 packages were installed in the conda base environment, so after running
I could run the python-GPIB scripts to interface with the Agilent.
Although, I did have to add a python2 kernel to jupyter/ipython, which I did in a separate conda environment:
conda create -n ipykernel_py2 python=2 ipykernel
source activate ipykernel_py2
python -m ipykernel install --user
I've installed pyepics on Donatella running
sudo yum install pyepics
Pip and ipython did not seem to be installed yet.
- There was a SR560+SR785 (not connected for measurement) placed near the X end which I moved; it is now behind the electronics rack by the X arm beam tube (~15m away).
- Also, for the AM measurement I moved the AG5395A from behind the PSL setup to the X end, where it now is.
- By toggling the XGREEN shutter, I noticed that the cavity was not resonant before I disconnected anything from the setup since the spot shape kept changing, but I proceeded anyway.
- Because Rana said that it was important for me to mention: the ~5 USD blue-yellow crocs (that I now use) work fine for me.
1. The cables were calibrated with the DC block in the A port (for a A/R measurement)
2. The cable to the PZT was disconnected from the pomona box and connected to the RF out of the NA, the PD output labelled 'GREEN_REFL' was also disconnected and connected to the B port via a DC block.
3. The ITMX was 'misaligned'. (This allowed the reflected green PD output as seen on the oscilloscope to stabilize.)
4. The PZT is modulated in frequency and the residual amplitude modulation (as observed in the measured reflected green light) is plotted, ref. Attachment 1. The parameters for the plotted data in the attachment were:
# AG4395A Measurement - Timestamp: Nov 07 2019 - 17:04:07
#---------- Measurement Parameters ------------
# Start Frequency (Hz): 10000.0, 10000.0
# Stop Frequency (Hz): 10000000.0, 10000000.0
# Frequency Points: 801, 801
# Measurement Format: LOGM, PHAS
# Measuremed Input: AR, AR
#---------- Analyzer Settings ----------
# Number of Averages: 8
# Auto Bandwidth: On, On
# IF Bandwidth: 300.0, 300.0
# Input Attenuators (R,A,B): 0dB 10dB 20dB
# Excitation amplitude = -10.0dBm
Update (19:13 7thNov19): When the ITMX was intentionally misaligned, Rana and I checked to see if the Oplevs were turned off and they were. But while I was casually checking the Oplevs again, they were on!
Not sure what to do about this or what caused it.
- At the X end, we set up the network analyzer to begin measurement of the AM transfer function by actuation of the laser PZT.
- The lid of the PDH optics setup was removed to make some checks and then replaced.
- From the PDH servo electronics setup the 'GREEN_REFL' and 'TO AUX-X LASER PZT' cables were removed for the measurement and then re-attached after.
- The signal today was too low to make a real measurement of the AM transfer function, but the GPIB scripts and interfacing was tested.
Gautam and I were talking about some modulation and demodulation and wondered what is the power combining situation for the triple resonant EOM installed 8 years ago. And we noticed that the current setup has additional ~5dB loss associated with the 3-to-1 power combiner. (Figure a)
N-to-1 broadband power combiners have an intrinsic loss of 10 log10(N). You can think about a reciprocal process (power splitting) (Figure b). The 2W input coming to the 2-port power splitter gives us two 1W outputs. The opposite process is power combining as shown in Figure c. This case, the two identical signals are the constructively added in the combiner, but the output is not 20Vpk but 14Vpk. Considering thge linearity, when one of the port is terminated, the output is going to be a half. So we expect 27dBm output for a 30dBm input (Figure d). This fact is frequently oversight particularly when one combines the signals at multiple frequencies (Figrue e). We can avoid this kind of loss by using a frequency-dependent power combiner like a diplexer or a triplexer.
[Mirko / Kiwamu]
The resonant box has been installed together with a 3 dB attenuator.
The demodulation phase of the MC lock was readjusted and the MC is now happily locked.
We needed more modulation depth on each modulation frequency and so for the reason we installed the resonant box to amplify the signal levels.
Since the resonant box isn't impedance matched well, the box creates some amount of the RF reflections (#5339).
In order to reduce somewhat of the RF reflection we decided to put a 3 dB attenuator in between the generation box and the resonant box.
(what we did)
+ attached the resonant box directly to the EOM input with a short SMA connector.
+ put stacked black plates underneath the resonant box to support the wight of the box and to relief the strain on the cable between the EOM and the box.
+ put a 3 dB attenuator just after the RF power combiner to reduce RF reflections.
+ readjusted the demodulation phase of the MC lock.
(Adjustment of MC demodulation phase)
The demodulation phase was readjusted by adding more cable length in the local oscillator line.
After some iterations an additional cable length of about 30 cm was inserted to maximize the Q-phase signal.
So for the MC lock we are using the Q signal, which is the same as it had been before.
Before the installation of the resonant box, the amplitude of the MC PDH signal was measured in the demodulation board's monitor pins.
The amplitude was about 500 mV in peak-peak (see the attached pictures of the I-Q projection in an oscilloscope). Then after the installation the amplitude decreased to 400 mV in peak-peak.
Therefore the amplitude of the PDH signal decreased by 20 %, which is not as bad as I expected since the previous measurement indicated 40 % reduction (#2586).
Koji and I taked about cleaning up some of the flaky cable situation on the PSL table a while ago. The changes were implemented and are documented in Attachment #1. Now the Pomona box between the Thorlabs HV Driver and the NPRO head is sitting on the PSL table (sandwiched between some teflon pieces I found in cabinet S4 along the south arm), and the cables between these two devices are better strain relieved. I turned off the Thorlabs HV supply while working on the PMC table. The IMC could be locked after this work. Probably won't solve the long standing FSS mysteries but probably can't hurt.
Unrelated to this work: I also removed a Bias tee that was just hanging out on top of the FSS electronics, which was used for the modeSpec project.
Here is a comparison of the response of various DoFs in our various RFPD sensors for two different CARM offsets. Even in the case of the smaller CARM offset of ~1kHz, we are several linewidths away from the resonance. Need to do some finesse modeling to make any meaningful statement about this - why is the CARM response in REFL11 apparently smaller for the smaller CARM offset?
If you mistrust my signal processing, the GPS times for which I ran the sensing lines are:
CARM offset = ~30kHz (arm transmission <0.02) --- 1257064777+5min
CARM offset = ~1kHz (arm transmission ~5) --- 1257065566+5min
There seems to be stronger-than-expected coupling between CARM and the 3f sensors.
A coarse calibration of the CARM error point (when on ALS control) is 7.040 +/- 0.030 kHz/ct. This corresponds to approximately 0.95nm/ct. I typically lose the PRMI lock when the CARM offset is ~0.2 cts, which means I am about 1kHz away from the resonance. This is >10 CARM linewidths.
The calibration was done by sweeping the CARM offset (no PRM) and identifying the arm cavity FSRs by looking for peaks in TRX / TRY. Attachment #1 shows the scan, while Attachment #2 shows a linear fit to the FSRs. In Attachment #2, the frequency axis is taken from the phase tracker servo, which was calibrated by injecting a "known" frequency with the Marconi, and there is good agreement to the expected FSR with 37.79 m long arm cavities. There is much more info in the scan (e.g. modulation depths, mode matching to the arm cavities etc) which I will extract later, but if anyone wants the data (pre-downsampled by me to have a managable filesize), it's attached as a .zip file in Attachment #3.
Full analysis tomorrow, but I collected sensing matrix measurements with lines driven in PRCL,MICH and CARM at a couple of CARM offsets. I also wanted to calibrate the CARM offset to physical units so I ran some scans of the CARM offset and collected the data so I can use the arm cavity FSR to calibrate CARM. Koji suggested using REFL165_I for PRCL and REFL165_Q for MICH control - this would allow us to see if the problem was with the 1f sideband only. While the lock could be established, we still couldn't push the arm powers above 10 without breaking the PRMI lock. While changing the CARM offset, we saw a significant shift in the DC offset level of the out-of-loop REFL33_I signal. Need to think about what this means...
I removed the Trillium T240 DAQ interface unit from 1X4 for investigation.
It was returned to the electronics rack and all the connections were re-made. Some details:
Update 445pm: Seems to have done something good - the old feedforward filters reduce the YAW RMS motion by a factor of a few. Pitch performance is not so good, maybe the filter needs re-training, but I see coherence, see Attachment #2 for the frequency domain WF.
I am still unable to achieve arm powers greater than TRX/TRY ~10 while keeping the PRMI locked. A couple of times, I was able to get TRY ~50, but TRX stayed at ~10, or even dropped a little, suggestive of a DARM offset? On the positive side, the ALS system seems to work pretty reliably, and I can keep the arms controlled by ALS for several tens of minutes.
I created a spreadsheet (Attached) by taking Koji's c1psl sheet from slow_channel_list and filtering out the channels that do not need an Acromag. I added in the QPD channels that are relevant to the PSL from the c1iool0 sheet.
I began mapping the PSL related Eurocrates connectors to their respective VME channels starting with the PMC electronics.
I am confused about the TTFSS interface (D040423): While it is a Eurocrate card, in the schematics it seems to have 50 pin connectors.
I found old wiring schematics that might help with identifying the channels once the connector issue is clarified.
I did some re-alignment of the POP beam on the IX in air table. Here are the details:
Tangentially related to this work - I took the nuclear option and did a hard reboot of the c1susaux Acromag crate on Sunday to fix the EPICS issue - it seems to be gone for now, see Attachment #5.
There are many versions of the POP22 signal path I found on the elog, e.g. this thread. But what I saw at the LSC rack was not quite in agreement with any of those. So here is the latest greatest version.
Since the 2f signals are mainly indicators of power buildups and are used for triggering various PDH loops, I don't know how critical some of these things are, but here are some remarks:
Thanks. Please update this wiki page too.
I reproduced Gautam's sketch of the 1x1 and 1x2 Eurocrates into a pdf image that contains links to the appropriate DCCs in the legend (see attachement).
PMC got unlocked at ~4am. I re-locked it. Also tweaked the input pointing into the cavity. The misalignment was mostly in pitch.
There was also a loud buzzing in the control room due to the audio cable being improperly seated in the mixer. I re-seated it.
Sigh... hard loch
While I was trying to lock the PRMI this evening, I noticed that I couldn't move the REFL beamspot on the CCD field of view by adjusting the slow bias voltages to the PRM. Other suspensions controlled by c1susaux seem to respond okay so at first glance it isn't a problem with the Acromag. Looking at the OSEM sensor input levels, I noticed that UL is much lower than the others - see Attachment #1, seems to have happened ~100 days ago. I plugged the tester box in to check if the problem is with the electronics or if this is an actual shorting of some pins on the physical OSEM as we had in the past. So PRM watchdog is shutdown for now and there is no control of the optic available as the cables are detached. I will replace the connections later in the evening.
Since I couldn't find anything wrong, I plugged the suspension back in - and voila, the suspect UL PD voltage level came back to a level consistent with the others! See Attachment #2.
Anyway, I had some hours of data with the tester box plugged in - see Attachment #3 for a comparison of the shadow sensor readout with the tester box (all black traces) vs with the suspension plugged in, local damping loops active (coloured traces). The sensing noise re-injection will depend on the specifics of the local damping loop shapes but I suspect it will limit feedforward subtraction possibilities at low frequencies.
However, I continue to have problems aligning the optic using the slow bias sliders (but the fast ones work just fine) - problem seems to be EPICS related. In Attachment #4, I show that even though I change the soft PITCH bias voltage adjust channel for the PRM, the linked channels which control the actual voltages to the coils take several seconds to show any response, and do so asynchronously. I tried restarting the modbus process on c1susaux, but the problem persists. Perhaps it needs a reboot of the computer and/or the acromag chassis? I note that the same problem exists for the BS and PRM suspensions, but not for ITMX or ITMY (didn't check the IMC optics). Perhaps a particular Acromag DAC unit is faulty / has issues with the internal subnet?
The POP beam coming out of the vacuum chamber is split by a 50/50 BS and half is diverted to the POP22/POP110/POPDC photodiode (Thorlabs PDA10CF) and the other half goes to the POP QPD. This optical layout is still pretty accurate. I looked at the data of the POPDC and POP QPD SUM channels while the dither alignment was running, to see if I could figure out what's up with the weird correlated dip in REFLDC and POPDC. While the POPDC channel shows some degradation as the REFLDC level goes down (=alignment gets better), the QPD sum channel shows the expected light level increase. So it could yet be some weird clipping somewhere in the beampath - perhaps at the 50/50 BS? I will lock the PRMI (no arms) and check...
I changed the shape of the low pass filter to reduce high frequency sensor noise injection into the MICH control signal. The loop stability isn't adversely affected (lost ~5 degrees of phase margin but still have ~50 degrees), while the control signal RMS is reduced by ~x10. This test was done with the PRMI locked on the carrier, need to confirm that this works with the arms controlled on ALS and PRMI lcoked on sideband.
Tried a bunch of things tonight.
One possibility is that the arm buildup is exerting some torque on the ITMs, which can also change the PRC cavity axis - as the buildup increases, the dominant component of the circulating field in the PRC comes from the leakage from the overcoupled arm cavity. We used to DC couple the ITM Oplev servos when locking the PRMI. The TRX level of 1 corresponds to ~5W of circulating power in the arm cavity, and the static radiation pressure force due to this circulating power is ~30 nN, rising up to 300nN as the TRX level hits 10. So for 1mm offset of the spot position on the ITM, we'd still only exert 300 pN m of torque. I don't see any transient in the Oplev error signals when locking the arm cavity as usual with POX/POY, but on timescales of several seconds, the Oplev error point shows ~3-5 urad of variation.
I set up a photodiode (PDA10CF) in the IFO REFL beampath and the Agilent NA is sitting on the east side of the PSL enclosure. This was meant to be just a first look, maybe the PDA10CF isn't suitable for this measurement. The measurement condition was - PRM aligned so we have a REFL beam (DC level = 8.4V measured with High-Z). Both ITMs and ETMs were macroscopically misaligned so that there isn't any cavity effects to consider. I collected noise around 11 and 55 MHz, and also a dark measurement, plots to follow. The optics were re-aligned to the nominal config but I left the NA on the east side of the PSL enclosure for now, in anticipation of us maybe wanting to tune something while minimizing a peak.
Attachment #1: Results of a coarse sweep from 5 MHz to 100 MHz. The broadband RIN level is not resolvable above the dark noise of the photodiode, but the peaks at the modulation frequencies (11 MHz, 55 MHz and 29.5 MHz) are clearly visible. Not sure what is the peak at ~44 MHz or 66 MHz. Come to think of it, why is the 29.5 MHz peak so prominent? The IMC cavity pole is ~4kHz so shouldn't the 29.5 MHz be attenuated by 80dB in transmission through the cavity?
Attachment #2: Zoomed in spectra with finer IF bandwidth around the RF modualtion frequencies. From this first measurement, it seems like the RIN/rad level is ~10^5, which I vaguely remember from discussions being the level which is best achieved in practise in the 40m in the past.
Check the RAM due to the EOM? Perhaps the pointing / polarization control into the EOM got degraded.
I looked at some signals for a 10 second period when the PRMI was locked with at some CARM offset, just before the PRMI lost lock, to see if there are any clues. I don't see any obvious signatures in this set of signals - if anything, the PRM is picking up some pitch offset, this is seen both at the Oplev error point and also in the POP QPD spot position. But why should this be happening as I reduce the CARM offset? The arm transmission is only ~5, so it's hard to imagine that the radiation pressure is somehow torquing the PRM. There are no angular feedback loops actuating on the PRM in this state except the local damping and Oplev loops.
The 1f signals are also changing their mean DC offset values, which may be a signature of a changing offset in the 3f MICH and PRCL error points? The MICH error signal is pretty noisy (maybe I can turn on some LPF to clean this up a bit), but I don't see any DC drift in the PRCL control signal.
I ran some sensing measurements at various CARM offsets to check if the PRCL-->REFL33 and MICH-->REFL165 signals were being rotated out of the sensing quadrature as I lowered the CARM offset - there was no evidence of this happening. See Attachment #2. Other possibilities:
The IMC went into some crazy state so I'm calling it for the night, need to think about what could be happening and take a closer look at more signals during the CARM offset reduction period for some clues...
In calculating the above numbers, I assumed a DC transimpedance of 10 khhms and an RF Transimpedance of ~800 V/A.
[Elog14480]: per these calculations, with the NewFocus 1611 PDs, we cannot achieve shot noise limited sensing for any power below the rated maximum for linear operation (i.e. 1mW). Moreover, the noise figure of the RF amplifier we use to amplify the sensed beat note before driving the delay-line frequency discriminator is unlikely to be the limiting noise source in the current configuration. Rana suggested that we get two Gain Blocks. These can handle input powers up to ~10dBm while still giving us plenty of power to drive the delay line. This way, we can (i) not compromise on the sacred optical gain, (ii) be well below the 1dB compression point (i.e. avoid nonlinear noise effects) and (iii) achieve a better frequency discriminant.
Temporary fix: While the gain blocks arrive, I inserted a 10dB (3dB) attenuator between the PSL+EX (PSL+EY) photodiode RF output and the ZHL-3A amplifiers. This way, we are well below the 1dB compression point of said RF amplifiers, and also below the 1dB compression point of the on-board Teledyne AP1053 amplifiers on the demodulator boards we use.
Nest steps: Rana is getting in touch with Rich Abbott to find out if there is any data available on the noise performance of the post-mixer IF amplifier stage in the 0.1 -30 Hz range, where the voltage and current noise of the AD829 OpAmps could be limiting the DFD performance. But in the meantime, the ALS noise seems good again, and there is no evidence of the sort of CARM/DARM coupling that motivated this investigation in the first place. Managed to execute several IR-->ALS transitions tonight in the PRFPMI locking efforts (next elog).
No new Teledyne AP1053s were harmed in this process - I'll send the 5 units back to Rich tomorrow.